T-Mobile believes that the % of people who change cellular providers in a year has changed based on market demographics . 19% is the well accepted population proportion of people who change providers. T Mobile samples 80 people: 20 of them say they switched providers in 2018. T- Mobile set alpha=0.07. Calculate the p value for this scenario enter to 3 decimal places. Before you solve, what is the check for this type of problem and did it pass?
Answer #1

Null and alternative hypotheses are

Ho : p = 0.19 Vs H1 : p 0.19

Test statistic Z = (p^ - p)/Sqrt [p*q/n]

Where p^ = 20/80 = 0.25

Z = ( 0.25 - 0.19)/sqrt[0.19*0.81/80]

Z = 1.37

p-value for Z = 1.37 and two tailed test

p-value = 2 * P( Z > 1.37)

p-value = 0.171